OpenStudy (anonymous):

Find x if :log(b8)+log(bx^2) =log(bx)

OpenStudy (anonymous):

Use those logarithm rules, silly. Go and rewrite the left side as \[log_{b}(something)\] It'll be quite easy after that.

OpenStudy (anonymous):

8x^2=x

OpenStudy (anonymous):

Note: x cannot be zero, because log(0) doesn't exist for any base

OpenStudy (anonymous):

x(8x-1) =0 x=(1/8)

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

remember the logarithm rules? specifically this one: \[log_{b}x+log_{b}y=log_{b}xy\] Use that to change the left side. Do that first, then the next step.

OpenStudy (anonymous):

how do you go from 8x2=x to x=1/8???

OpenStudy (anonymous):

Chrissy, do what I said, you'll see. You are also reading elec's post wrongly.

OpenStudy (anonymous):

Are my posts all on one line? I think there is something wrong with this site. I usually space things ot but obviously you can't see the spaces I put in.

OpenStudy (anonymous):

I understand that logb8+logbx^2= log8x^2, what is the next step Ian??

OpenStudy (anonymous):

Oh, okay. \[log_{b}8+log_{b}x^{2}=log_{b}x\] \[log_{b}8x^{2}=log_{b}x\] \[8x^{2}=x\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Oh, do you want after all of that? \[8x^{2}-x=0\] This is sort of like a quadratic equation, if you think about it.

OpenStudy (anonymous):

I have not started quadratic equations yet

OpenStudy (anonymous):

Really? logarithms are usually after those. Factorize the left hand side.

OpenStudy (anonymous):

I can only prime factorize

OpenStudy (anonymous):

You cannot factorize algebraic expressions? For example: ab+ac factorized is a(b+c) The factors are a and (b+c)

OpenStudy (anonymous):

I can only prime factorize

OpenStudy (anonymous):

not yet no. in class, we have started linear equations, and what I would call simple logarithms. If I had known that I needed to know more stuff before starting these logarithms I would I waited.

OpenStudy (anonymous):

ohhh. Well, to explain it: \[8x^{2}-x=0\] you factorize it. Do you know the distributive property? You can use hat to see this step. \[x(8x-1)=0\] so something multiplied by something is 0. In order for that to be true, one of the somethings must be zero, or both of them are zero. so... x=0 or 8x-1=0 x cannot be 0 because the log of 0 is undefined. So it must be 8x-1=0 8x=1 x=1/8

OpenStudy (anonymous):

I can only prime factorize

OpenStudy (anonymous):

I dont get it, but Thanks for your help